OpenStudy (anonymous):

Jesus please help me. I am not getting any right answer and have been asking a lot How can one fifthx − 2 = one thirdx + 8 be set up as a system of equations? 5y − 5x = −10 3y − 3x = 24 5y − 5x = −10 3y + 3x = 24 5y + x = −10 3y + x = 24 5y − x = −10 3y − x = 24

OpenStudy (anonymous):

One fifth of x is \[\frac{ 1 }{ 5 } \times x\] one third x is \[\frac{ 1 }{ 3 } \times x\]

OpenStudy (anonymous):

But this is just one equation and not set of equations. Please see your question again.

OpenStudy (anonymous):

I know how its written It's asking how to set up as system of equation and it does not look anything like those

OpenStudy (anonymous):

y = 1/5x-2 and y = 1/3x + 8 so you see my problem